# Weighted algebraic topology, II

### 05 May 2026 v1

@@ 5.2. The valuation of a path
More generally, we consider now a set-theoretical path $a\colon\mathbb{I}\to X$ (a mapping of sets) in a $\mu$-space. (A restriction to Lipschitz paths $\delta\mathbb{I}\to X$ would have no relevance here.)
(a) We define the valuation $v(a)$, also written as $v_{X}(a)$ or $v_{\mu}(a)$, by the following functions
(75) $v_{\underline{t}}(a)=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\mu(a(t_{j-1}),a(t_{j})),\qquad v(a)={\rm sup}_{\underline{t}}\,v_{\underline{t}}(a)\in{\bf w},$
@@
We always have $v(a)\geqslant\mu(a(0),a(1))$. If $X$ is a $\delta$-metric space, $v(a)\geqslant 0$ is also called the length of $a$ and written $L(a)$, as in [G3], Definition 1.2, and in [G4], 6.1.8.
@@
(76) $\begin{array}[]{c}v^{+}(a)=L_{X^{+}}(a)={\rm sup}_{\underline{t}}\,\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\mu^{+}(a(t_{j-1}),a(t_{j}))\in{\bf w}^{+},\\[5.0pt] v^{+}(a)\,\geqslant\,v(a).\end{array}$
@@
(c) Absolute value. If $\mu$ is linear, its absolute value $|\mu|(x,y)=|\mu(x,y)|$ is easily seen to be a finite symmetric $\delta$-metric (generally called a pseudometric in the literature).
The total variation of the path $a$ will be its $|\mu|$-valuation
(79) $\begin{array}[]{l}L_{|\mu|}(a)={\rm sup}_{\underline{x}}\,\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,|\mu|(x_{j-1},x_{j})={\rm sup}_{\underline{x}}\,\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,|(\varphi(x_{j})-\varphi(x_{j-1})|\\[5.0pt] \qquad\;\;\;\;=v^{+}_{\mu}(a)+v^{-}_{\mu}(a).\end{array}$
@@ 5.3. Lemma
If $\mu$ is linear, with potential $\varphi$, for any path $a\colon\mathbb{I}\to X$ (with reversed path $a^{\sharp}$):
(80) $\begin{array}[]{c}v(a)=\mu(a(0),a(1))=\varphi(a(1))-\varphi(a(0))\in\mathbb{R},\\[3.0pt] v(a^{\sharp})=-v(a),\qquad\qquad v^{-}(a)=v^{+}(a^{\sharp}).\end{array}$
@@
The first relation is obvious and the second is a consequence. As to to the third, we recall that $\mu(y,x)=-\mu(x,y)$ and note that, for $\lambda\in\mathbb{R}$, we have $\lambda=\lambda^{+}-(-\lambda)^{+}$.
Letting $\underline{t}$} be a finite increasing sequence of points in $\mathbb{I}$, with $0=t_{0}\leqslant t_{1}\leqslant...\leqslant t_{n}=1$ and $p_{j}=a(t_{j})$, we have by linearity the following relations
$\begin{array}[]{c}v(a)=\mu(p_{0},p_{1})=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\mu(p_{j-1},p_{j})=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\mu^{+}(p_{j-1},p_{j})\!-\!\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\mu^{+}(p_{j},p_{j-1}),\\[5.0pt] \raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\mu^{+}(p_{j-1},p_{j})=v(a)+\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\mu^{+}(p_{j},p_{j-1}),\\[5.0pt] v^{+}(a)=v(a)+{\rm sup}_{\underline{t}}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\mu^{+}(p_{j},p_{j-1})))=v(a)+v^{+}(a^{\sharp}),\end{array}$
@@ 5.4. Examples and remarks
(b) For a path $a$ in the linear $\mu$-space $\mu\mathbb{R}$:
@@
- also the weight $a||$ (of $a\colon\delta\mathbb{I}\to\mu\mathbb{R}$) can be infinite: a classical example is $a(t)=\sqrt{t}$,
- for an increasing path, $v^{+}(a)=v(a)=a(1)-a(0)$ and $v^{-}(a)=0$,
- for a decreasing path , $v^{-}(a)=-v(a)=a(0)-a(1)$ and $v^{+}(a)=0$.
(c) For a path $a$ in $\delta_{0}\mathbb{R}$, $L(a)$ is the upper valuation of $a$ in $\mu\mathbb{R}$ and can be infinite. For an increasing (resp. decreasing) path, L(a)=a(1)a(0)L(a)=a(1)-a(0) (resp. $L(a)=0$).
(d) For a (set-theoretical) path $a$ in a discrete $\mu$-space $\mu_{d}X=\delta_{d}X$, $L(a)$ is 0 or $\infty$, according to the fact that $a$ is constant or not.
In a chaotic $\mu$-space $\mu_{c}X$, $v(a)={\overline{\infty}}$. In a chaotic $\delta$-space $\delta_{c}X$, $L(a)=0$.
@@ 5.5. Theorem (Path valuation)
Let $X$ be a $\mu$-metric space.
The valuation of paths satisfies the following properties, where $a\colon\mathbb{I}\to X$ is a set-theoretical path, $e_{x}$ is the constant path at $x\in X$, aba*b is the concatenation of two consecutive paths, $\rho\colon\mathbb{I}\to\mathbb{I}$ is a partial reparametrisation (an increasing map) and $a||\leqslant\infty$ is the Lipschitz weight of $a\colon\mathbb{I}\to X$ as a mapping defined on the $\delta$-space $\delta\mathbb{I}$.
@@
Point (e) follows from the fact that we are computing the mapping $v$ on the same sequences: if $(t_{j})$ is an increasing sequence in $\mathbb{I}$, $(\rho(t_{j}))$ is also; conversely, given $(t_{j})$, there is an increasing sequence $(s_{j})$ such that $\rho(s_{j})=t_{j}$, for all indices $j$. (Note also that $\rho(0)=0$ and $\rho(1)=1$.)
@@
Finally, points (f) and (g) are obvious, because:
@@ 5.6. Spaces with distinguished paths
This can be extended to $\mu$-spaces (still using the reflective topology): the axioms of $\delta$-spaces are still satisfied (by Theorem 5.5(a), (b), (d)), and $\mu$-maps preserve the directed paths (by 5.5(g)).
Yet, we have already seen that this extension cannot be satisfactory: $\mu\mathbb{R}$ and several related $\mu$-spaces would get the chaotic topology. Using the final topology on $X$ for the affordable paths might be preferable, but is not an extension of the positive case, and would give strange topologies, of little interest, to basic spaces like all $\delta\mathbb{R}^{n}$, for $n>1$.
@@
Finally, this is another evidence of the vague relationship of relative metrics to topology. As claimed in the Introduction, Part III will give a better solution to all these problems.
@@ 6. Tensor products, other examples
The category $\mu\mathsf{Mtr}$ has a ‘natural’ symmetric monoidal closed structure, that extends the structure of $\delta\mathsf{Mtr}$ [Lw, G3, G4], with the same formulas. We also examine the relationship between the tensor product and the cartesian product. The elementary example of the Introduction is reviewed in 6.5; the gravitational and elastic potential in 6.6, 6.7.
@@ 6.1. Tensor product
(a) The tensor product $X\otimes Y$ of $\mu$-spaces is the cartesian product of the underlying sets, endowed with the $l_{1}$-type $\mu$-metric (instead of the $l_{\infty}$-type $\mu$-metric of the categorical product)
(81) $\mu((x,y),(x^{\prime},y^{\prime}))=\mu_{X}(x,x^{\prime})+\mu_{Y}(y,y^{\prime}).$
The unit is the $\delta$-singleton $\delta\top$ (see 2.4(b)). On the other hand, $X\otimes\mu\top=\mu_{c}|X|$, for all affordable $\mu$-spaces $X$.
This construction solves the usual universal problem, with respect to mappings which are 1-Lipschitz in each variable. The exponential $Z^{Y}$ is the set of 1-Lipschitz maps $Y\to Z$ equipped with the $\mu$-metric
(82) $\mu(h,k)={\rm sup}_{y\in Y}\,\mu_{Z}(h(y),k(y))\qquad\;\;(h,k\colon Y\to Z\mbox{ in }\mu\mathsf{Mtr}).$
The proof of the adjunction is standard, and can be found in point (d). The tensor powers of the basic $\mu$-structures of the line are written as $\delta\mathbb{R}^{\otimes n}$, $\mu\mathbb{R}^{\otimes n}$ and $\delta_{0}\mathbb{R}^{\otimes n}$.
(b) Restricting to positive $\delta$-metrics, the cartesian and tensor product satisfy the following inequalities
@@
(c) For $\mu$-metrics we still have
(84) $X\otimes Y\leqslant 2(X\!\times\!Y)\qquad\qquad(\text{in }\mu\mathsf{Mtr}),$
but the first inequality in (83) fails (also for symmetric $\mu$-spaces). For instance, the $\mu$-spaces $\delta\top\!\times\!\mu\top=\delta\top$ and $\delta\top\otimes\mu\top=\mu\top$ give the reversed inequality, and are not even Lipschitz-isomorphic.
@@
(85) $f\colon X\otimes Y\to Z,\qquad g\colon X\to Z^{Y}\qquad\qquad(\text{in }\mu\mathsf{Mtr}),$
@@
Given $f$, the relation $\mu_{Z}(f(x,y),f(x^{\prime},y^{\prime}))\leqslant\mu_{X}(x,x^{\prime})+\mu_{Y}(y,y^{\prime})$ shows that:
- every map $g(x)=f(x,-)\colon Y\to Z$ is 1-Lipschitz, because (taking into account that $\mu_{X}(x,x)\leqslant 0$):
$\mu_{Z}(f(x,y),f(x,y^{\prime}))\leqslant\mu_{X}(x,x)+\mu_{Y}(y,y^{\prime})\leqslant\mu_{Y}(y,y^{\prime}),$
- the map $g\colon X\to Z^{Y}$ is 1-Lipschitz, because we have, in the same way:
$\begin{array}[]{c}\mu_{Z}(g(x)(y),g(x^{\prime})(y))=\mu_{Z}(f(x,y),f(x^{\prime},y))\\[3.0pt] \qquad\qquad\qquad\qquad\leqslant\;\mu_{X}(x,x^{\prime})+\mu_{Y}(y,y)\leqslant\mu_{X}(x,x^{\prime}).\end{array}$
@@ 6.2. Theorem (Properties of the tensor product)
Let $X,Y$ be $\mu$-metric spaces.
(a) The reflector $^{\,{\scriptstyle{\wedge}}\,}\colon\mu\mathsf{Mtr}\to\;!\mu\mathsf{Mtr}$ is a strict monoidal functor
@@
and forms, with the embedding $\mu\mathsf{Mtr}\subset\mu\mathsf{Mtr}$, a strict adjunction of monoidal categories.
(b) The coreflector $^{\,{\scriptstyle{\vee}}\,}\colon\mu\mathsf{Mtr}\to\;!\mu\mathsf{Mtr}$ is a lax monoidal functor
@@
(c) As to the reflective and coreflective preorders (in 4.3):
@@
We replace the pair $X,Y$ with a finite family of $\mu$-metric spaces $X_{1},...,X_{n}$; this will give a simpler notation and a clearer proof. An element of the set $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}_{i}|X_{i}|$ is written as ${\underline{x}}=(x_{i})$. We use the notation $\mu_{X}(x,x^{\prime})=X(x,x^{\prime})$, which comes from viewing a $\mu$-metric space as an enriched category (in 2.1).
(a) We now prove that $(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}_{i}\,X_{i})^{\,{\scriptstyle{\wedge}}\,}=\;\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}_{i}\,X_{i}^{\,{\scriptstyle{\wedge}}\,}$. The $\mu$-metric of the latter satisfies:
@@
Since $\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i}^{\,{\scriptstyle{\wedge}}\,}$ is symmetric, we have $\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i}^{\,{\scriptstyle{\wedge}}\,}\leqslant(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i})^{\,{\scriptstyle{\wedge}}\,}$. The reversed inequality is more subtle: take a sequence of $n+1$ points ${\underline{z}}^{j}$, which varies from $\underline{x}$} to $\underline{y}$} by changing one coordinate at a time
@@
and apply the triangular inequality:
@@
Now, ${\underline{z}}^{j-1}$ and ${\underline{z}}^{j}$ only differ at the $j$-th coordinate ($x_{j}$ or $y_{j}$, respectively); restricting the domain of the ‘inf’ in the right term above to those sequences in $(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i})^{\,{\scriptstyle{\wedge}}\,}$ where only the $j$-th coordinate changes, we get the $\delta$-metric $X_{j}^{\,{\scriptstyle{\wedge}}\,}$, and the inequality:

### 13 May 2026 v2

@@ 5.2. The valuation of a path
More generally, we consider now a set-theoretical path $a\colon\mathbb{I}\to X$ (a mapping of sets) in a $\rho$-space. (A restriction to Lipschitz paths $\delta\mathbb{I}\to X$ would have no relevance here.)
(a) We define the valuation $v(a)$, also written as $v_{X}(a)$ or $v_{\rho}(a)$, by the following functions:
(68) $v_{\underline{t}}(a)=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\rho(a(t_{j-1}),a(t_{j})),\qquad v(a)={\rm sup}_{\underline{t}}\,v_{\underline{t}}(a)\in{\bf w},$
@@
We always have $v(a)\geqslant\rho(a(0),a(1))$. If $X$ is a $\delta$-metric space, $v(a)\geqslant 0$ is also called the length of $a$ and written $L(a)$, as in [G3], Definition 1.2, and in [G4], 6.1.8.
@@
(69) $\begin{array}[]{c}v^{+}(a)=L_{X^{+}}(a)={\rm sup}_{\underline{t}}\,\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\rho^{+}(a(t_{j-1}),a(t_{j}))\in{\bf w}^{+},\\[5.0pt] v^{+}(a)\,\geqslant\,v(a).\end{array}$
@@
(c) Absolute value. If $\rho$ is linear, its absolute value $|\rho|(x,y)=|\rho(x,y)|$ is easily seen to be a finite symmetric $\delta$-metric (generally called a pseudometric in the literature).
The total variation of the path $a$ will be its $|\rho|$-valuation
(72) $\begin{array}[]{l}L_{|\rho|}(a)={\rm sup}_{\underline{x}}\,\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,|\rho|(x_{j-1},x_{j})={\rm sup}_{\underline{x}}\,\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,|(\Phi(x_{j})-\Phi(x_{j-1})|\\[5.0pt] \qquad\;\;\;\;=v^{+}_{\rho}(a)+v^{-}_{\rho}(a).\end{array}$
@@ 5.3. Lemma
If $\rho$ is linear, with potential $\Phi$, for any path $a\colon\mathbb{I}\to X$ (with reversed path $a^{\sharp}$)
(73) $\begin{array}[]{c}v(a)=\rho(a(0),a(1))=\Phi(a(1))-\Phi(a(0))\in\mathbb{R},\\[3.0pt] v(a^{\sharp})=-v(a),\qquad\qquad v^{-}(a)=v^{+}(a^{\sharp}).\end{array}$
@@
The first relation is obvious and the second is a consequence. As to to the third, we recall that $\rho(y,x)=-\rho(x,y)$ and note that, for $\lambda\in\mathbb{R}$, we have $\lambda=\lambda^{+}-(-\lambda)^{+}$.
Letting $\underline{t}$} be any finite increasing sequence of points in $\mathbb{I}$, with $0=t_{0}\leqslant t_{1}\leqslant...\leqslant t_{n}=1$ and $p_{j}=a(t_{j})$, we have by linearity the following relations:
$\begin{array}[]{c}v(a)=\rho(p_{0},p_{1})=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho(p_{j-1},p_{j})=\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho^{+}(p_{j-1},p_{j})\!-\!\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho^{+}(p_{j},p_{j-1}),\\[5.0pt] \raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho^{+}(p_{j-1},p_{j})=v(a)+\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho^{+}(p_{j},p_{j-1}),\\[5.0pt] v^{+}(a)=v(a)+{\rm sup}_{\underline{t}}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,\rho^{+}(p_{j},p_{j-1})))=v(a)+v^{+}(a^{\sharp}),\end{array}$
@@ 5.4. Examples and remarks
(b) For a path $a$ in the linear $\rho$-space $\rho\mathbb{R}$
@@
- also the weight $a||$ (of $a\colon\delta\mathbb{I}\to\rho\mathbb{R}$) can be infinite: a classical example is $a(t)=\sqrt{t}$,
- if $a$ is increasing, $v^{+}(a)=v(a)=a(1)-a(0)$ and $v^{-}(a)=0$,
- if $a$ is decreasing, $v^{-}(a)=-v(a)=a(0)-a(1)$ and $v^{+}(a)=0$.
(c) For a path $a$ in $\delta_{0}\mathbb{R}$, the length $L(a)$ is the upper valuation of $a$ in $\rho\mathbb{R}$ and can be infinite. For an increasing (resp. decreasing) path, L(a)=a(1)a(0)L(a)=a(1)-a(0) (resp. $L(a)=0$).
(d) For a (set-theoretical) path $a$ in a discrete $\rho$-space $\rho_{D}X=\delta_{D}X$, $L(a)$ is 0 or $\infty$, according to the fact that $a$ is constant or not.
In a chaotic $\rho$-space $\rho_{C}X$, $v(a)={\overline{\infty}}$. In a chaotic $\delta$-space $\delta_{C}X$, $L(a)=0$.
@@ 5.5. Theorem (Path valuation)
Let $X$ be a $\rho$-metric space.
The valuation of paths satisfies the following properties, where $a\colon\mathbb{I}\to X$ is a set-theoretical path, $e_{x}$ is the constant path at $x\in X$, aba*b is the concatenation of two consecutive paths, $\varphi\colon\mathbb{I}\to\mathbb{I}$ is a partial reparametrisation (an increasing map) and $a||\leqslant\infty$ is the Lipschitz weight of $a\colon\mathbb{I}\to X$ as a mapping defined on the $\delta$-space $\delta\mathbb{I}$.
@@
Point (e) follows from the fact that we are computing the mapping $v$ on the same sequences: if $(t_{j})$ is an increasing sequence in $\mathbb{I}$, $(\varphi(t_{j}))$ is also; conversely, given $(t_{j})$, there is an increasing sequence $(s_{j})$ such that $\varphi(s_{j})=t_{j}$, for all indices $j$. (Note also that $\varphi(0)=0$ and $\varphi(1)=1$.)
@@
Finally, points (f) and (g) are obvious, because
@@ 5.6. Spaces with distinguished paths
This can be extended to $\rho$-spaces (still using the reflective topology): the axioms of $\delta$-spaces are still satisfied (by Theorem 5.5(a), (b), (d)), and $\rho$-maps preserve the directed paths (by 5.5(g)).
Yet, we have already seen that this extension cannot be satisfactory: $\rho\mathbb{R}$ and several related $\rho$-spaces would get the chaotic topology. Using the final topology on $X$ for the affordable paths might be preferable, but is not an extension of the positive case, and would give strange topologies, of little interest, to basic spaces like all $\delta\mathbb{R}^{n}$, for $n>1$.
@@
Finally, this is another evidence of the vague relationship of real metrics to topology. As claimed in the Introduction, Part III will give a better solution to all these problems.
@@ 6. Tensor products, other examples
The category $\rho\mathsf{Mtr}$ has a ‘natural’ symmetric monoidal closed structure, that extends the structure of $\delta\mathsf{Mtr}$ [L1, G3, G4], with the same formulas. We also examine the relationship between the tensor product and the cartesian product. The elementary example of the Introduction is reviewed in 6.5; the gravitational and elastic potential in 6.6, 6.7.
@@ 6.1. Tensor product
(a) The tensor product $X\otimes Y$ of $\rho$-spaces is the cartesian product of the underlying sets, endowed with the $l_{1}$-type $\rho$-metric (instead of the $l_{\infty}$-type $\rho$-metric of the categorical product)
(74) $\rho((x,y),(x^{\prime},y^{\prime}))=\rho_{X}(x,x^{\prime})+\rho_{Y}(y,y^{\prime}).$
The unit is the $\delta$-singleton $\delta\top$ (see 2.4(b)). On the other hand, $X\otimes\rho\top=\rho_{C}|X|$, for all affordable $\rho$-spaces $X$.
This construction solves the usual universal problem, with respect to mappings which are 1-Lipschitz in each variable. The exponential $Z^{Y}$ is the set of 1-Lipschitz maps $Y\to Z$ equipped with the $\rho$-metric
(75) $\rho(h,k)={\rm sup}_{y\in Y}\,\rho_{Z}(h(y),k(y))\qquad\;\;(h,k\colon Y\to Z\mbox{ in }\rho\mathsf{Mtr}).$
The proof of the adjunction is standard, and can be found in point (d). The tensor powers of the basic $\rho$-structures of the line are written as $\delta\mathbb{R}^{\otimes n}$, $\rho\mathbb{R}^{\otimes n}$ and $\delta_{0}\mathbb{R}^{\otimes n}$.
(b) Restricting to positive $\delta$-metrics, the cartesian and tensor product satisfy the following inequalities:
@@
(c) For $\rho$-metrics we still have
(77) $X\otimes Y\leqslant 2(X\!\times\!Y)\qquad\qquad(\text{in }\rho\mathsf{Mtr}),$
but the first inequality in (76) fails (also for symmetric $\rho$-spaces). For instance, the $\rho$-spaces $\delta\top\!\times\!\rho\top=\delta\top$ and $\delta\top\otimes\rho\top=\rho\top$ give the reverse inequality, and are not even Lipschitz-isomorphic.
@@
(78) $f\colon X\otimes Y\to Z,\qquad g\colon X\to Z^{Y}\qquad\qquad(\text{in }\rho\mathsf{Mtr}),$
@@
Given $f$, the relation $\rho_{Z}(f(x,y),f(x^{\prime},y^{\prime}))\leqslant\rho_{X}(x,x^{\prime})+\rho_{Y}(y,y^{\prime})$ shows that
- every map $g(x)=f(x,-)\colon Y\to Z$ is 1-Lipschitz, because (taking into account that $\rho_{X}(x,x)\leqslant 0$)
$\rho_{Z}(f(x,y),f(x,y^{\prime}))\leqslant\rho_{X}(x,x)+\rho_{Y}(y,y^{\prime})\leqslant\rho_{Y}(y,y^{\prime}),$
- the map $g\colon X\to Z^{Y}$ is 1-Lipschitz, because we have, in the same way
$\begin{array}[]{c}\rho_{Z}(g(x)(y),g(x^{\prime})(y))=\rho_{Z}(f(x,y),f(x^{\prime},y))\\[3.0pt] \qquad\qquad\qquad\qquad\leqslant\;\rho_{X}(x,x^{\prime})+\rho_{Y}(y,y)\leqslant\rho_{X}(x,x^{\prime}).\end{array}$
@@ 6.2. Theorem (Properties of the tensor product)
Let $X,Y$ be $\rho$-metric spaces.
(a) The reflector $^{\,{\scriptstyle{\wedge}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr}$ is a strict monoidal functor
@@
and forms, with the embedding $\rho\mathsf{Mtr}\subset\rho\mathsf{Mtr}$, a strict adjunction of monoidal categories.
(b) The coreflector $^{\,{\scriptstyle{\vee}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr}$ is a lax monoidal functor
@@
(c) As to the reflective and coreflective preorders (in 4.3)
@@
We replace the pair $X,Y$ with a finite family of $\rho$-metric spaces $X_{1},...,X_{n}$; this will give a simpler notation and a clearer proof. An element of the set $\raisebox{1.93747pt}{${\mbox{$\prod$}}$}_{i}|X_{i}|$ is written as ${\underline{x}}=(x_{i})$. We use the notation $\rho_{X}(x,x^{\prime})=X(x,x^{\prime})$, which comes from viewing a $\rho$-metric space as an enriched category (in 2.1).
(a) We now prove that $(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}_{i}\,X_{i})^{\,{\scriptstyle{\wedge}}\,}=\;\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}_{i}\,X_{i}^{\,{\scriptstyle{\wedge}}\,}$. The $\rho$-metric of the latter satisfies
@@
Since $\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i}^{\,{\scriptstyle{\wedge}}\,}$ is symmetric, we have $\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i}^{\,{\scriptstyle{\wedge}}\,}\leqslant(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i})^{\,{\scriptstyle{\wedge}}\,}$. The reverse inequality is more subtle: take a sequence of $n+1$ points ${\underline{z}}^{j}$, which varies from $\underline{x}$} to $\underline{y}$} by changing one coordinate at a time
@@
and apply the triangular inequality
@@
Now, ${\underline{z}}^{j-1}$ and ${\underline{z}}^{j}$ only differ at the $j$-th coordinate ($x_{j}$ or $y_{j}$, respectively); restricting the domain of the ‘inf’ in the right term above to those sequences in $(\raisebox{1.80832pt}{${\mbox{$\bigotimes$}}$}\,X_{i})^{\,{\scriptstyle{\wedge}}\,}$ where only the $j$-th coordinate changes, we get the $\delta$-metric $X_{j}^{\,{\scriptstyle{\wedge}}\,}$, and the inequality